OpenStudy (anonymous):

Find the sum of all positive rational numbers that are less than 5, and that have a denominator of 30 when written in lowest terms.

OpenStudy (anonymous):

I have no idea?

OpenStudy (debbieg):

Hmmm... well, to be a positive, rational number that is less than 5 and has a den'r of 30, the number has to be of the form \[\frac{ a }{ 30 }\]where some very specific limitations can be placed on a. Can you think of what they are?

OpenStudy (anonymous):

they have to have no common factors?

OpenStudy (debbieg):

Hmm.... wait. This is more complicated than I initially thought, lol. But yes, they have to have no common factors... so the numerator can't be a multiple of 2, 3, or 5.

OpenStudy (debbieg):

We also know a lower and upper bound on a.... right?

OpenStudy (anonymous):

0-5?

OpenStudy (debbieg):

no. first of all for lower bound, we are looking for positive number so must have a>0, so a=1 is the smallest possible a.

OpenStudy (debbieg):

And since we want a/30<5, we need to have a<.... what?

OpenStudy (anonymous):

150

OpenStudy (debbieg):

Right... so 0<a<150, and no factors of 2,3 or 5. Now I'm kind of stuck. Sorry, I thought I had an idea where to go, but I'm not sure what to do next. It actually wouldn't be that hard to cook up the list by hand, but I doubt that such "brute force" is what's expected. There must be some slicker method.

OpenStudy (debbieg):

Wait, I have an idea.... remember, what we need is a way to sum up these terms in the numerator. Once we have that sum, they are all over the LCD of 30. So if we take\[\large \sum_{n=1}^{149}n\]that gives ALL the numbers in that interval. the problem with that, of course, is that it includes all the numbers that have factors in common with 30. But what if we do:\[\large \sum_{n=1}^{149}n-\sum_{n=1}^{74}2n-\sum_{n=1}^{49}3n-\sum_{n=1}^{29}5n\]Do you see what that does? Now if I could just remember how to evaluate those sums... lol.

OpenStudy (debbieg):

OK, I think the sums are n(n+1)/2, but we have to pull out the constants. Annnnd... I'm talking to myself so I'm outta here. Respond and/or tag me if you want further help.

OpenStudy (anonymous):

Ughh...I'm so rude. I had to get off because of an...emergency. Sorry!

OpenStudy (debbieg):

lol 'sok, thanks for coming back to explain. You are allowed to have a life. :)

OpenStudy (anonymous):

First, let's look at all such rational numbers between 0 and 1.

OpenStudy (anonymous):

100

OpenStudy (debbieg):

What?? @jonjenkins7653 the rational numbers are all supposed to be less than 5. Did you read my discussion above?

OpenStudy (anonymous):

Aww...thanks for understanding debbie

OpenStudy (debbieg):

I think it works.\[\large \sum_{n=1}^{149}n-\sum_{n=1}^{74}2n-\sum_{n=1}^{49}3n-\sum_{n=1}^{29}5n\]\[=\large \sum_{n=1}^{149}n-2\sum_{n=1}^{74}n-3\sum_{n=1}^{49}n-5\sum_{n=1}^{29}n\]\[=\frac{ 149(150) }{2 }-2\frac{ 74(75) }{2 }-3\frac{ 49(50) }{2 }-5\frac{ 29(30) }{2 }\]

OpenStudy (anonymous):

@DebbieG Now you have to add back on the multiples of 2 and 3 (i.e. 6) because you have taken them away twice. Similarly with the multiples of 10 and 15. And finally take away the multiples of 30, since you took those away three times but added them back on three times!

OpenStudy (debbieg):

I think?? That would give the numerator..... Oh wow! Yes of course, I see what you mean @Erin001001 !

OpenStudy (anonymous):

this is a great question!

OpenStudy (anonymous):

wait...what? I'm so confused

OpenStudy (anonymous):

sum all the primes from 7 to 150 and divide by 30. any other number will be a multiple of 2, 3, 5 or some combination of thos numbers.

OpenStudy (anonymous):

of course, 150 isn't prime but i hope you get the idea...

OpenStudy (debbieg):

I was originally thinking of just doing something with primes. Wasn't sure if that was the way to go though.

OpenStudy (anonymous):

the numerator must be relatively prime to 2, 3 & 5... so all the primes and multiples that don't include those. so sorry, not just the primes...

OpenStudy (debbieg):

That's basically what I was getting at above, when I said doing it by "brute force". lol That would work, since the number of terms in the sum is reasonably small. But it seems like there should be a more algorithmic way to approach it, lol. :)

OpenStudy (anonymous):

the sum of the terms is 2266

OpenStudy (debbieg):

Right @pgpilot326 ... I think if done by brute force, you just have to take all integers from 1 to 149 that aren't a multiple of 2,3 or 5. So a number like 77 is not prime, but IS part of the sum.

OpenStudy (anonymous):

@orple8 2266? Is that 2266/30? Intuitively, 2266 seems too big to me.

OpenStudy (anonymous):

yeah, i jumped to soon... it just seems that the including everything then adding and subtracting different multiples is a bit confusing.